Calculate goodwill, if it is to be calculate at 3years’ purchase of the super profit; The

firm started business with a capital of Rs.4,00,000. The normal rate of earning in this class of

business is 15%. The firm earned Rs.74,000 as profit during the year.


1
Expert's answer
2021-09-09T18:10:51-0400

average profit= Rs 74000

normal profit"=400000\times 15/100=Rs60000"

super profit= average profit -normal profit=74000-60000=Rs14000

good will = super profit "\times" no. of years purchase "=14000\times 3=Rs42000"